OpenStudy (anonymous):

8.6 plus 8.6 times a number equals the sum of the number and 69.8 what is the number

OpenStudy (anonymous):

something like 4.05 \[(8.6+8.6)x=69.8\] \[17.2 x=69.8\] \[17.2 x/17.2=69.8/17.2=x\]

OpenStudy (anonymous):

did this help u